# Channel-wise Parallelizable Spiking Neuron with Multiplication-free Dynamics and Large Temporal Receptive Fields

### v1

@@ 6 Conclusion
In this paper, we introduce a novel parallelizable spiking neuron model named mul-free channel-wise PSN, which employs the channel-wise convolutions to process the input sequences, avoids the large neuron order by sawtooth dilations, and gets rid of floating multiplications by efficient bit shift operations. The considerations of accelerating the training of SNNs with the proposed neuron models are also discussed in detail. Experimental results demonstrate that mul-free channel-wise PSN achieves significant performance improvements in temporal classification tasks, showcasing its superior capability to capture long-term dependencies. Our methods solve the dilemma of performance and computational costs of spiking neuron models, and our acceleration methods will benefit future research as a practical reference.
@@
A    Acceleration Details of Spiking Neuron Layer
A.1    Time-first/last + Vmap + MM
@@
$\bm{A}[:][i][j]=\begin{cases}\begin{aligned} W_{q}[:][k-1-\frac{i-j}{d}],\ &i-d(k-1)\leq j\leq i\\ &\&\ (i-j)\%d=0\end{aligned}\\ 0,\text{otherwise}\end{cases},$ (19)
where [:]delimited-[]:[:][ : ] means the slice operation.
@@
$\bm{A}[:][j][i]=\begin{cases}\begin{aligned} W_{q}[:][k-1-\frac{i-j}{d}],\ &i-d(k-1)\leq j\leq i\\ &\&\ (i-j)\%d=0\end{aligned}\\ 0,\text{otherwise}\end{cases}.$ (20)
@@
.0.1 A.2    Time-first/last + Custom CUDA Kernel
@@
where pad represents padding $k-1$ zeros on the left of $\bm{X}$ over the time dimension $T$, and $\star$ denotes the convolution operation on the $T$ dimension of $X$ using $W$.
@@
Beyond PyTorch (cuDNN), a custom CUDA implementation for two data layouts is also considered. To avoid the reshape and incident memory copying, we design CUDA kernels by OpenAI Triton for processing both data layouts directly. Specifically, we manually implement the Eqs.(22)-(25) using the Triton framework. We design a custom autograd function, where the aforementioned kernel functions are called in the forward and backward methods. Noting that the convolution operation in Eq.(23) is consistent with that in Eq.(22), so the triton kernel function remains the same. Taking the time-first layout as an example, Eq.(22) and (24) could be implemented as Algorithms 2 and 3, respectively.
@@
Input: The input sequence pointer $X_{ptr}$, weight matrix pointer $W_{ptr}$, the output sequence pointer $H_{ptr}$, point to the first address of tensors, shaped ad $[T+k-1,N,C,H,W]$, $[C,k]$ and $[T,N,C,H,W]$, respectively.
[…]
@@
Input: The grad of output pointer $O_{ptr}$, the input sequence pointer $X_{ptr}$, the grad of weight pointer $W_{ptr}$, point to the first address of tensors, shaped ad $[T,N,C,H,W]$, $[T+k-1,N,C,H,W]$ and $[C,k]$, respectively.
[…]
@@
B    Acceleration Details of Stateless Layer
Dataset Optimizer Batch Size Epoch Learning Rate Scheduler
Sequential CIFAR10/100 AdamW 128 256 0.001 CosineAnnealingLR
Pixel CIFAR10 AdamW 128 128 0.001 CosineAnnealingLR
SHD Adam (wd=1e-5) 256 150 0.001 for weights 0.1 for delay CosineAnnealingLR for weights OneCycleLR for delay
DVS-Lip Adam (wd=1e-4) 32 200 fixed 3e-4 for 0-100 epochs (1e-4, 5e-6) for 100-200 epochs CosineAnnealingLR
Table 4: Training hyper-parameters for different datasets.
Stateless layers include the convolutional, batch normalization, pooling, and linear layers. When using the time-first data layout, the stateless layers can be accelerated by fusing the time dimension and the batch dimension in SpikingJelly. More specifically, the data layout changes as $(T,N,*)\rightleftharpoons(TN,*)$ before and after processing of the stateless layers. Then GPUs regard the time-step dimension as the batch dimension, leading to fully parallel computing over time-steps. It is worth noting that the dimension fusion is nearly no cost because the time and batch dimensions are physically adjacent in memory. The reshape operation only changes the view of tensors and does not involve the memory copying.
When using the time-last layout, the dimension fusion method of SpikingJelly cannot be applied expect for the batch normalization layer, which only requires that the channel dimension should be the 1-th dimension. Both layouts can be satisfied by reshape without additional memory R/W. For the convolutional and pooling layer, we introduce two new methods, the vectorizing map provided in PyTorch and the high-dimension convolution/pooling that has been used in the Lava framework, a software framework for neuromorphic computing.
The vectorizing map vectorizes the stateless layers to process the input sequence with the $(N,...,T)$ layout over the last dimension $T$, then the computation over time-steps is in parallel. This method actually implies a reshape operation $(N,...,T)\rightleftharpoons(T,N,...)$ when spliting and concatenating the sequence. The high-dimension convolution/pooling use the $(n+1)-$D convolution/pooling to implement the $n$-D convolution/pooling with weight as 1 and stride as 1 in the additional dimension, which is similar to using 2-D convolution to implement Eq.(17) discussed before. This method is also in parallel, while the main drawback is that the high-dimension convolution/pooling is complex and not as efficient as the dimension fusion method [Fang et al., 2023a].
C    Neuron Quantization
To alleviate the internal covariate shift along the temporal and batch dimension, increase the numerical stability of the model, we use the batch normalization to implement the learnable threshold $V_{th}$. Eq.(18) could be rewrite as:
@@ C    Neuron Quantization
where $\bm{\gamma}\in\mathbb{R}^{C}$ and $\bm{\beta}\in\mathbb{R}^{C}$ are the learnable weight, initialized as $1$ and $-1$. $\bm{\mu}_{\mathcal{B}}\in\mathbb{R}^{C}$ and $\bm{\sigma}^{2}_{\mathcal{B}}\in\mathbb{R}^{C}$ are the mean and variance of the input over the dimension $C$. Specifically, at train time, they are the mean and biased variance of the input sequence; at inference time, they are the moving average of the mean and unbiased variance of the input sequence on the training stage, which means $\bm{\mu}_{\mathcal{B}}$ and $\bm{\sigma}^{2}_{\mathcal{B}}$ are invariant during inference.
Since our quantization goal is to use the efficient bit shifting operation to replace the multiplication, the convolution layer and the batch normalization layer could be fused to reduce computation during inference, so we need to quantize the fused weights during training. The formula for the fusing of convolution and batch normalization can be represented as follows:
@@
D    Network Structure
@@
Table 5: Network structure for different datasets.
Table 5 illustrates the details of the network structure for different datasets. c128k3s1 represents convolution layer with output channels $=128$, kernel size $=3$ and stride $=1$, BN is the batch normalization. SN is the mul-free channel-wise PSN, APk2s2 is the avg-pooling layer with kernel size $=2$ and stride $=2$, FC256 represents the fully connected layer with output feature $=256$. RB128 is the residual block with output channels $=128$, Dcls256 is the dilated convolution with learnable spacings with output channels $=256$, DP is the dropout layer. LIF(Vth=1e9) represents an LIF spiking neuron the threshold $=1e9$, and the membrane potential is the output, which could be thought of the moving average of the input current. 3D_c64k577s122p233 represents the 3D convolution layer with output channels $=64$, kernel size $=(5,7,7)$, stride $=(1,2,2)$ and padding $=(2,3,3)$. AAPk1 is the adaptive avg-pooling layer with output feature $=1$. Stateful FC 100 is a FC layer with stateful synapses.
E    Setting of Experiments
@@
Sequential CIFAR The data augmentation techniques include random mixup with $p=1$ and $\alpha=0.2$, random cutmix with $p=1$ and $\alpha=1$, random choice between the two mix methods with $p=0.5$, random horizontal flip with $p=0.5$, trivial augmentation, normalization, random erasing with $p=0.1$, and label smoothing with the amount $0.1$ [Fang et al., 2023b]. The number of channels is $128$. The surrogate function is the arctan surrogate function $\sigma(x)=\frac{\alpha}{2(1+(\frac{\pi}{2}\alpha x)^{2})}$ with $\alpha=2$.
Pixel CIFAR All is the same as Sequential CIFAR.
SHD No specific augmentation method is implemented. The surrogate function is also the arctan surrogate function with $\alpha=5$.
DVS-Lip The data augmentation techniques include center cropped size $=96\times 96$, then random cropped size $=88\times 88$, random horizontal flip with $p=0.5$, 2D spatial mask with mask num $=4$ and maximum length$=20$, random choice between zoom in and zoom out with $p=0.5$ and max scale $=26$, temporal mask with mask num $=6$ and maximum length $=18$ [Dampfhoffer et al., 2024]. The surrogate function is $\sigma(x)=\frac{1}{1+\alpha x^{2}}$ with $\alpha=10$.
F    Experimental Results
The original data for Figure 6 and Figure 7 are shown in Table 6 and Table 7, respectively. Notably, the training duration of our neuron is with the quantization operation, i.e. perform the convolution operation twice, so it is inherently slower than PSN on the training stage.
Order Dataset sequential CIFAR100 pixel CIFAR 10 (w/o dilation) pixel CIFAR 10 (w/ dilation)
@@ F    Experimental Results
Table 6: Original test accuracy (%) of Figure 6.
Method T 2 4 8 16 32
PSN/SlidingPSN 17.86 23.04 27.47 44.18 88.12
Time_first(k=2) 28.23 42.10 493.01 623.1 881.87
Time_first(k=4) 43.63 771.39 888.15 1154.12
Time_first(k=8) 1379.37 1413.37 1696.01
Time_last(k=2) 30.36 42.43 286.33 369.94 547.64
Time_last(k=4) 52.44 493.25 622.84 845.38
Time_last(k=8) 1379.70 1422.11 1657.56
Autoselect(k=2) 15.06 25.33 45.82 84.91 181.06
Autoselect(k=4) 24.67 45.59 77.60 231.65
Autoselect(k=8) 46.40 78.22 193.85
Table 7: Original training duration (s/epoch) of Figure 7.
In Figure 4, we mention that the gradient of Eq.(15) is jump points, which is detrimental to the network. On the sequence CIFAR dataset, we find that the network is still able to learn quite well. However, on the DVS-Lip dataset, as shown in Figure 8, using the original ste gradient causes the network to crash, resulting in the training and testing accuracy suddenly dropping to $1\%$. The reason is that the gradients appear to be the Not a Number (NaN) values.

### v2

@@ 6 Conclusion
In this paper, we introduce a novel parallelizable spiking neuron model named mul-free channel-wise PSN, which employs the channel-wise convolutions to process the input sequences, avoids the large neuron order by sawtooth dilations, and gets rid of floating multiplications by efficient bit-shift operations. The considerations of accelerating the training of SNNs with the proposed neuron models are also discussed in detail. Experimental results demonstrate that mul-free channel-wise PSN achieves significant performance improvements in temporal classification tasks, showcasing its superior capability to capture long-term dependencies. Our methods solve the dilemma of performance and computational costs of spiking neuron models, and our acceleration methods will benefit future research as a practical reference.
Acknowledgments and Disclosure of Funding
This work is supported by National Science and Technology Innovation 2030 Major Project (No. 2025ZD0215501), Guangdong S&T Programme 2024B0101010003 and National Natural Science Foundation of China (62236009).
@@
Appendix A Acceleration Details of Spiking Neuron Layer
A.1 Acceleration Methods
Based on the aforementioned background, we design the following accelerating implementations as candidates:
(1) Time-last + Vmap + Conv1d: this method uses the vectorizing map function (Vmap) in PyTorch to vectorize Conv1d to process the input sequence with the (N,C,,T)(N,C,*,T) layout over the "*" dimension. The reshape operations $(N,C,...,T)\rightleftharpoons(N,C,*,T)$ are nearly cost-free because the reshaped dimensions are adjacent physically.
(2) Time-last + Conv2d: this method is similar to the implementation (1), but processes the input sequence with the (N,C,,T)(N,C,*,T) by Conv2d and sets the weight and stride in the "*" dimension as 1, rather than by Vmap.
(3) Time-first/last + Vmap + MM: this method uses Vmap to vectorize matrix multiplication (MM) to process inputs over channels ($c$ in Eq.(17)). Refer to Appendix A.2 for more details about how the weights for MM are generated.
(4) Time-first/last + Custom CUDA Kernel: this method avoids reshape operations and can be used for any memory layout. However, the convolutions in PyTorch are highly optimized, i.e., implemented by the official NVIDIA CUDA Deep Neural Network (cuDNN) library, which might be much faster than custom implementations. Refer to Appendix A.3 for more details.
Note that if the spiking neuron layer implementations adopt the time-last layout, the stateless layers should also use the same layout. Otherwise, reshape operations between time-first and time-last layouts will cause great latency. Correspondingly, the time batch dimension fusion method to accelerate stateless layers in SpikingJelly cannot be applied. In Appendix B, we fully discuss the acceleration method of the stateless layer for the time-last layout.
A.2 Details of Time-first/last + Vmap + MM
@@
$\bm{A}[:][i][j]=\begin{cases}W_{q}[:][k-1-\frac{i-j}{d}],\ &i-d(k-1)\leq j\leq i\ \ \&\ \ (i-j)\%d=0\\ 0,\text{otherwise}\end{cases},$ (19)
where $[:]$ means the slice operation.
@@
$\bm{A}[:][j][i]=\begin{cases}W_{q}[:][k-1-\frac{i-j}{d}],\ &i-d(k-1)\leq j\leq i\ \ \&\ \ (i-j)\%d=0\\ 0,\text{otherwise}\end{cases}.$ (20)
@@
A.3 Details of Time-first/last + Custom CUDA Kernel
@@
where pad represents padding k1k-1 zeros on the left of $\bm{X}$ over the time dimension $T$, and $\star$ denotes the convolution operation on the $T$ dimension of $X$ using $W$.
@@
Beyond PyTorch (cuDNN), a custom CUDA implementation for two data layouts is also considered. To avoid the reshape and incident memory copying, we design CUDA kernels by OpenAI Triton for processing both data layouts directly. Specifically, we manually implement the Eqs.(22)-(25) using the Triton framework. We design a custom autograd function, where the aforementioned kernel functions are called in the forward and backward methods. Note that the convolution operation in Eq.(23) is consistent with that in Eq.(22), so the triton kernel function remains the same. Taking the time-first layout as an example, Eq.(22) and (24) could be implemented as Algorithms 2 and 3, respectively.
@@
Input: The input sequence pointer $X_{ptr}$, weight matrix pointer $W_{ptr}$, the output sequence pointer $H_{ptr}$, point to the first address of tensors, shaped ad [T+k1,N,C,H,W][T+k-1,N,C,H,W], $[C,k]$ and $[T,N,C,H,W]$, respectively.
1: Utilize the triton autotune method to determine the $\text{BLOCK\_SIZE\_NHW}(BN)$ and $\text{BLOCK\_SIZE\_C}(BC$)
@@
Input: The grad of output pointer $O_{ptr}$, the input sequence pointer $X_{ptr}$, the grad of weight pointer $W_{ptr}$, point to the first address of tensors, shaped ad $[T,N,C,H,W]$, [T+k1,N,C,H,W][T+k-1,N,C,H,W] and $[C,k]$, respectively.
1: Utilize the triton autotune method to determine the $\text{BLOCK\_SIZE\_NHW}(BN)$ and $\text{BLOCK\_SIZE\_C}(BC)$
@@
Appendix B Acceleration Details of Stateless Layer
Stateless layers include the convolutional, batch normalization, pooling, and linear layers. When using the time-first data layout, the stateless layers can be accelerated by fusing the time dimension and the batch dimension in SpikingJelly. More specifically, the data layout changes as $(T,N,*)\rightleftharpoons(TN,*)$ before and after processing of the stateless layers. Then GPUs regard the time-step dimension as the batch dimension, leading to fully parallel computing over time-steps. It is worth noting that the dimension fusion is nearly no cost because the time and batch dimensions are physically adjacent in memory. The reshape operation only changes the view of tensors and does not involve memory copying.
When using the time-last layout, the dimension fusion method of SpikingJelly cannot be applied except for the batch normalization layer, which only requires that the channel dimension be the 1-th dimension. Both layouts can be satisfied by a reshape without additional memory R/W. For the convolutional and pooling layer, we introduce two new methods: the vectorizing map provided in PyTorch and the high-dimension convolution/pooling that has been used in the Lava framework, a software framework for neuromorphic computing.
[…]
Appendix C Neuron Quantization
To reduce the internal covariate shift along the temporal and batch dimensions, and increase the numerical stability of the model, we use batch normalization to implement the learnable threshold $V_{th}$. Eq.(18) could be rewrite as:
@@ Appendix C Neuron Quantization
where $\bm{\gamma}\in\mathbb{R}^{C}$ and $\bm{\beta}\in\mathbb{R}^{C}$ are the learnable weight, initialized as $1$ and 1-1. $\bm{\mu}_{\mathcal{B}}\in\mathbb{R}^{C}$ and $\bm{\sigma}^{2}_{\mathcal{B}}\in\mathbb{R}^{C}$ are the mean and variance of the input over the dimension $C$. Specifically, at train time, they are the mean and biased variance of the input sequence; at inference time, they are the moving average of the mean and unbiased variance of the input sequence on the training stage, which means $\bm{\mu}_{\mathcal{B}}$ and $\bm{\sigma}^{2}_{\mathcal{B}}$ are invariant during inference.
Since our quantization goal is to use the efficient bit-shift operation to replace the multiplication, the convolution layer and the batch normalization layer could be fused to reduce computation during inference, so we need to quantize the fused weights during training. The formula for the fusion of convolution and batch normalization can be represented as follows:
@@
Appendix D Network Structure
Table 7 illustrates the details of the network structure for different datasets. c128k3s1 represents convolution layer with output channels =128=128, kernel size =3=3 and stride =1=1, BN is the batch normalization. SN is the mul-free channel-wise PSN, APk2s2 is the avg-pooling layer with kernel size =2=2 and stride =2=2, FC256 represents the fully connected layer with output feature =256=256. RB128 is the residual block with output channels =128=128, Dcls256 is the dilated convolution with learnable spacings with output channels =256=256, DP is the dropout layer. LIF(Vth=1e9) represents a LIF spiking neuron the threshold =1e9=1e9, and the membrane potential is the output, which could be thought of the moving average of the input current. 3D_c64k577s122p233 represents the 3D convolution layer with output channels =64=64, kernel size =(5,7,7)=(5,7,7), stride =(1,2,2)=(1,2,2) and padding =(2,3,3)=(2,3,3). AAPk1 is the adaptive avg-pooling layer with output feature =1=1. Stateful FC 100 is an FC layer with stateful synapses.
Table 7: Network structure for different datasets.
@@ Appendix D Network Structure
Appendix E Setting of Experiments
@@
Sequential CIFAR The data augmentation techniques include random mixup with $p=1$ and $\alpha=0.2$, random cutmix with $p=1$ and $\alpha=1$, random choice between the two mix methods with $p=0.5$, random horizontal flip with $p=0.5$, trivial augmentation, normalization, random erasing with $p=0.1$, and label smoothing with the amount $0.1$ [15]. The number of channels is $128$. The surrogate function is the arctan surrogate function $\sigma(x)=\frac{\alpha}{2(1+(\frac{\pi}{2}\alpha x)^{2})}$ with $\alpha=2$.
Pixel CIFAR All parameters and experimental settings are the same as Sequential CIFAR.
SHD Augmentation methods include spatio jitter with var $0.55$, uniform noise with number $n=35$, drop event with $p=0.05$, drop event chunk with $p=0.3$ and max drop chunk length $l=0.02$. The surrogate function is also the arctan surrogate function with $\alpha=5$.
DVS-Lip The data augmentation techniques include center cropped size $96\times 96$, then random cropped size $88\times 88$, random horizontal flip with $p=0.5$, 2D spatial mask with mask num =4=4 and maximum length=20=20, random choice between zoom in and zoom out with $p=0.5$ and max scale =26=26, temporal mask with mask num =6=6 and maximum length =18=18 [42]. The surrogate function is $\sigma(x)=\frac{1}{1+\alpha x^{2}}$ with $\alpha=10$.
Table 8: Training hyper-parameters for different datasets.
Dataset Optimizer Batch Size Epoch Learning Rate Scheduler
Sequential CIFAR10/100 AdamW 128 256 0.001 CosineAnnealingLR
Pixel CIFAR10 AdamW 128 128 0.001 CosineAnnealingLR
SHD
Adam
(wd=1e-5)
128 150
5e-4 for weights
0.05 for delay
CosineAnnealingLR for weights
OneCycleLR for delay
DVS-Lip
Adam
(wd=1e-4)
32 200
fixed 3e-4 for 0-100 epochs
(1e-4, 5e-6) for 100-200 epochs
CosineAnnealingLR
Appendix F Details of the Energy Estimation
By combining the neuron operation counts in Table 1 with the corresponding per-operation energy consumption, the theoretical neuronal energy consumption can be estimated, as shown in Table 9. When the neuron order $k$ reaches its maximum value $T$, the maximum theoretical energy consumption of our neuron is $0.515T^{2}+1.415T$, which is approximately $9\times$ lower than that of PSN for large $T$. This result clearly demonstrates the superior energy efficiency of our approach in hardware-constrained scenarios.
Table 9: Operations and energy of different spiking neurons during inference.
Neuron Operations Energy (pJ)
PSN $(T^{2}+T)\times$ ADD, $T^{2}\times$ MUL $4.6T^{2}+0.9T$
Ours $((T+\frac{1-k}{2})\cdot k+T)\times$ ADD, $((T+\frac{1-k}{2})\cdot k)\times$ SHIFT $1.03(T+\frac{1-k}{2})\cdot k+0.9T$
[…]
$E_{\text{Synaptic}}$ $=E_{MAC}\times\text{FL}_{\text{SNN Conv}}^{1}+E_{AC}\times\left(\sum_{n=2}^{N}\text{SOP}^{n}_{\text{SNN Conv}}+\sum_{m=1}^{M}{\text{SOP}^{m}_{\text{SNN FC}}}\right)$ (29)
$\text{SOPs}(l)$ $=fr\times T\times\text{FLOPs}(l)$ (30)
Appendix G Experimental Results
Table 10 presents the original data depicted in Figure 5, illustrating the performance of mul-free channel-wise PSN across varying orders and datasets.
Table 10: Test accuracy (%) of the multiplication-free channel-wise PSN, corresponding to the data presented in Figure 5.
Order Dataset
Sequential
CIFAR100
Pixel CIFAR 10
(w/o dilation)
Pixel CIFAR 10
(w/ dilation)
@@ Appendix G Experimental Results
In addition, we report extra results on the H20 GPU and the A100 GPU with different batch sizes in Tab.11. The results and conclusions are consistent with using the A100 GPU and batch size 128. Notably, the training duration of our neuron is with the quantization operation, i.e., perform the convolution operation twice, so it is inherently slower than PSN on the training stage.
Table 11: Training durations (s/epoch).
Method $T=2$ $T=4$ $T=8$ $T=16$ $T=32$
GPU=H20, batch size = 128
PSN 10.07 16.79 27.47 50.27 102.42
Autoselect ($k=2$) 12.63 22.25 40.99 84.65 153.95
Autoselect ($k=4$) 23.09 42.85 83.22 154.06
Autoselect ($k=8$) 45.33 81.44 154.44
GPU=A100, batch size = 32
PSN 20.11 23.57 32.88 50.13 88.62
Autoselect ($k=2$) 29.83 39.73 56.40 105.82 215.23
Autoselect ($k=4$) 35.11 50.91 104.40 225.88
Autoselect ($k=8$) 52.17 96.41 187.18
GPU=A100, batch size = 64
Autoselect 13.72 17.46 27.47 45.84 85.79
Autoselect ($k=2$) 19.75 31.94 47.13 96.79 197.29
Autoselect ($k=4$) 32.98 47.55 101.88 217.38
Autoselect ($k=8$) 48.86 105.86 245.72
For the training speed across different neuron implementations (PSN, vanilla, and autoselect), each method is trained for two epochs, and the training durations of the second epoch are set as the results. In this way, the results have accounted for statistical validity and warm-up of GPUs and eliminated the influence of model loading and other initialization processes. Hence, the reported speed is the speed at the steady state.
@@
In Figure 3(b), we mention that the gradient of Eq.(15) has jump points, which is detrimental to the network. On the sequence CIFAR dataset, we find that the network is still able to learn quite well. However, on the DVS-Lip dataset, as shown in Figure 7, using the original ste gradient causes the network to crash, resulting in the training and testing accuracy suddenly dropping to $1\%$. The reason is that the gradients appear to be the Not a Number (NaN) values.
Appendix H Limitations
Although we have designed multiple acceleration methods that significantly improve the training speed of the vanilla mul-free channel-wise PSN, its training speed is still slightly slower than that of PSN. This is mainly due to the additional overhead induced by quantization, and the memory read/write operations caused by Vmap during input/output processing in our SNNs being slower compared to dimension fusion. Nonetheless, the speed gap is not significant.